OpenStudy (abbles):

Where does the pi/2 comes from?

OpenStudy (abbles):

Where do they get the pi/2 after the cos? I must be missing something

OpenStudy (phi):

good question. They pulled it out of their hat (or at least, did not explain how they got it) you can find the slope from 0,0 to 1,1, i..e 1 and the slope of the line from 0,0 to -3,3 i.e. -1 the slopes are "negative reciprocals" \[ m_1 = - \frac{1}{m_2} \] and that means the lines are perpendicular, and form a 90º angle, or \( \frac{\pi}{2}\) radians

OpenStudy (abbles):

Ah, interesting. Because for the question that comes right after the lesson goes like this... The angle between vectors (2, 3) and (-3, 2) is _ degrees. Would I calculate the slope to find the angle?

OpenStudy (phi):

no. we use the two different ways to compute the dot product the "easy" way (at least for me) is multiply corresponding components and add the other way is v dot w = |v| |w| cos x we know v dot w (from the first way) we know the lengths (or magnitudes) of v and w that means we can "solve for cos x"

OpenStudy (abbles):

The dot product will be zero, right? so then I plug that into this: v dot w = |v| |w| cos x what do the v and w values represent?

OpenStudy (phi):

I just mean two random vectors v and w the | v | means the length of v (the square root of the sum of the squares of the components) here you know the dot product is 0 so you have 0 = |(2, 3)| | (-3, 2) | cos x we could figure out the actual lengths (both are sqr(8+4)= sqr(12)= 2sqr(3) ) but you will end up with cos x = 0 and \[ x = \cos^{-1}0 = 90º\]

OpenStudy (phi):

** we could figure out the actual lengths (both are sqr(9+4)= sqr(13)

OpenStudy (abbles):

|(2, 3)| | (-3, 2) | does this mean we add the vectors?

OpenStudy (phi):

no. the | v | means "find the length of vector v" for example, for the vector (1,1,) we could do this: 

OpenStudy (phi):

for the vector (2,3) (2 over, 3 up) we use c^2= a^2 + b^2 where a is 2, b is 3 c^2 = 2^2 + 3^2 c^2= 4+9 c^2= 13 c= sqr(13) that is using Pythagorean Theorem but people just square the components: (2,3) --> do 2*2 + 3*3 = 4+9 = 13 then take the square root : sqr(13) that is the "length" of the vector.

OpenStudy (phi):

the formula \[ v \cdot w =\ |v|\ |w| \cos \theta\] means multiply the length of vector v times the length of vector w then multiply by the angle between the two vectors (assuming you know it)

OpenStudy (abbles):

Oh! So my equation would look like this: 0 = 2sqrt13costheta And solve for theta?